OpenStudy (anonymous):

A particle moving in a straight line has a velocity given by v(t)=2sin(t)-1. What is the rate of change of velocity, called the average acceleration, of the particle over the interval (π/6) ≤ t ≤ (π/4)?

OpenStudy (anonymous):

do you know how to find the derivative of that?

OpenStudy (anonymous):

one second

OpenStudy (anonymous):

that's what they are asking you to do.

OpenStudy (anonymous):

the derivative of velocity makes your acceleration function

OpenStudy (anonymous):

so is it 2cos(t)-1?

OpenStudy (anonymous):

one second

OpenStudy (anonymous):

i could be wrong, but i think it is just cos(t)

OpenStudy (asnaseer):

I believe the derivative would you the "instantaneous" acceleration. what you are being asked to find is the "average" acceleration.

OpenStudy (anonymous):

or 2cos(t)

OpenStudy (asnaseer):

average acceleration = (final_velocity - initial_velocity) / time_taken

OpenStudy (anonymous):

so i dont need to find the derivative asnaseer? and ya i was wrong, the derivative would be 2cos(t)

OpenStudy (anonymous):

would the derivative over the interval give you the same answer?

OpenStudy (asnaseer):

bleh123: thats right - no need to find derivatives

OpenStudy (anonymous):

so what do i do with the interval it gives?

OpenStudy (anonymous):

he doesn't have time though.

OpenStudy (anonymous):

*she

OpenStudy (asnaseer):

time_taken = start_time - end_time

OpenStudy (asnaseer):

sorry - other way around

OpenStudy (anonymous):

ah ok

OpenStudy (asnaseer):

start_time = \(\pi/6\) end_time = \(\pi/4\)

OpenStudy (anonymous):

umm so how would i solve this?

OpenStudy (asnaseer):

I have given you all the information you need to work it out

OpenStudy (anonymous):

would initial velocity be 0?

OpenStudy (asnaseer):

yes - that looks right

OpenStudy (anonymous):

the equation is pretty ugly.

OpenStudy (anonymous):

2sin(t)-1 over pi/6 -pi/4

OpenStudy (asnaseer):

initial_velocity = \(v(\pi/6)\) final_velocity = \(v(\pi/4)\)

OpenStudy (anonymous):

-_- so i set it up wrong?

OpenStudy (anonymous):